OpenStudy (anonymous):

Multiply: x^2-9/3x * 5x/x+3

OpenStudy (anonymous):

\[\frac{ x^2-9 }{ 3x }* \frac{ 5x }{ x+3 }\]

OpenStudy (anonymous):

answers: \[\frac{ 5(x-3) }{ 3 }\] \[\frac{ 5(x+3) }{ 3 }\] \[\frac{ (x-3) }{ 3 }\] \[\frac{ (x+3) }{ 3 }\]

OpenStudy (anonymous):

@Directrix

Directrix (directrix):